OpenStudy (calebxjr):

omg pls help ill give fan medal Whats 9+10=

candycove (candycove):

9 + 10 = kittens.

OpenStudy (calebxjr):

ok thank you

candycove (candycove):

You're welcome.

OpenStudy (calebxjr):

lol

OpenStudy (faiqraees):

No no 9+10 equals 19 kittens

OpenStudy (calebxjr):

o i thought it was 11kittens

OpenStudy (lord2clash):

9+10= 19 19/2 = 9.5 9.5 + fish = Kittens sorry i thought i might elaborate on how candy got his answer

OpenStudy (calebxjr):

thanks for the answers ill fan

candycove (candycove):

rip